8. Will a precipitate form if 10.0mL of a 2.00x10-5M solution of AgNO3 (aq) is combined with 10.0mL of a 0.0340 M K2CrO4 (aq)? The Ksp for Ag2CrO4 is 1.1x10-12.
Answer #1

Lets find the concentration after mixing for AgNO3

Concentration after mixing = mol of component / (total volume)

M(AgNO3) after mixing = M(AgNO3)*V(AgNO3)/(total volume)

M(AgNO3) after mixing = 2.0E-5 M*10.0 mL/(10.0+10.0)mL

M(AgNO3) after mixing = 1*10^-5 M

Lets find the concentration after mixing for K2CrO4

Concentration after mixing = mol of component / (total volume)

M(K2CrO4) after mixing = M(K2CrO4)*V(K2CrO4)/(total volume)

M(K2CrO4) after mixing = 0.034 M*10.0 mL/(10.0+10.0)mL

M(K2CrO4) after mixing = 1.7*10^-2 M

So, we have now

[Ag+] = 1*10^-5 M

[CrO42-] = 1.7*10^-2 M

At equilibrium:

Ag2CrO4 <----> 2 Ag+ + CrO42-

Qsp = [Ag+]^2[CrO42-]

Qsp = (1*10^-5)^2*(1.7*10^-2)

Qsp = 1.7*10^-12

we have,

Ksp = 1.1*10^-12

Since Qsp is greater than ksp, precipitate will form

Answer: yes, precipitate will form